### v2

# Prospects for the Detection of High-Redshift Gamma-Ray Bursts in the Era of EP and SVOM

Abstract
Gamma-ray bursts (GRBs) are a promising probe of the high-redshift Universe, but their detection remains observationally challenging. In this work, we explore the detectability of high-$z$ GRBs by the Wide-field X-ray Telescope (WXT) aboard the Einstein Probe (EP) and the coded-mask gamma-ray imager (ECLAIRs) aboard the Space-based multi-band astronomical Variable Objects Monitor (SVOM). Using a population synthesis model calibrated to $Swift$ GRB observations, we develop a tool to estimate high-$z$ GRB detection rates for instruments with specific energy bands and sensitivities. Our results indicate that EP/WXT could detect $\sim 5.1^{+3.4}_{-2.4}$ (with 68% confidence level) GRBs annually at $z>6$, compared to $\sim 0.7^{+1.0}_{-0.4}$ $\mathrm{events\,yr^{-1}}$ at $z>6$ for SVOM/ECLAIRs. While EP cannot independently determine redshifts (requiring optical/near-infrared follow-up), its assumed $\sim 30\%$ follow-up efficiency yields $\sim 1.5^{+1.0}_{-0.7}$ confirmed $z>6$ GRBs annually. SVOM, equipped with dedicated follow-up telescopes, will promptly identify high-$z$ candidates deserving deep near-infrared spectroscopy to ensure robust confirmation of high-$z$ GRBs. We anticipate that EP and SVOM will open new avenues for utilizing enlarged samples of high-$z$ GRBs to explore the early Universe. Moreover, EP will assemble a substantial sample of soft, low-luminosity GRBs at low-to-intermediate redshifts, providing critical insights into the structure of GRB jets.
Gamma-ray bursts (629) — Star formation (1569) — Observational cosmology (1146)
1 Introduction
Gamma-ray bursts (GRBs), among the most energetic explosive transients in our Universe, are detectable up to very high redshifts. The most distant spectroscopically confirmed burst to date is GRB 090423 at $z=8.2$ (Salvaterra et al., 2009b; Tanvir et al., 2009). As such, GRBs represent a promising probe, complementary to quasar and galaxy surveys, for exploring the early Universe (see Salvaterra 2015; Wang et al. 2015 for reviews). In particular, long GRBs with durations $T_{90}>2\,\mathrm{s}$ (where $T_{90}$ is defined as the time interval containing 90% of the prompt emission; Kouveliotou et al. 1993) arise from the core collapse of massive stars [^1] Notably, two nearby long-duration bursts, GRB 211211A and GRB 230307A, exhibit kilonova signatures indicative of compact binary mergers, implying that at least some long GRBs originate from non-collapsar progenitors (Rastinejad et al., 2022; Troja et al., 2022; Yang et al., 2022; Dichiara et al., 2023; Levan et al., 2024). Given that merging compact binary progenitors have significantly longer average lifetimes than collapsars, the fraction of merger-origin long-duration GRBs (especially those bright enough to be detectable) is expected to decrease substantially at high redshifts compared to $z=0$. (e.g., Woosley 1993; Paczyński 1998; Woosley & Bloom 2006), establishing them as powerful tracers of cosmic star formation (e.g., Totani 1997; Chary et al. 2007; Yüksel et al. 2008; Kistler et al. 2009; Wang & Dai 2009; Robertson & Ellis 2012; Wang 2013; Wei et al. 2014; Matsumoto et al. 2024). Moreover, GRB afterglows are expected to be observable up to $z\sim 20$ (Ciardi & Loeb, 2000; Gou et al., 2004), enabling spectroscopic afterglow studies to provide independent constraints on the reionization history, ionizing photon escape fraction, and cosmic metal enrichment history, and $\mathrm{21\,cm}$ absorption features (Ioka & Mészáros, 2005; Totani et al., 2006, 2014; Wang et al., 2012; Chornock et al., 2013, 2014; Ciardi et al., 2015; Hartoog et al., 2015; Tanvir et al., 2019; Saccardi et al., 2023; Fausey et al., 2025a, b; Šoltinský et al., 2025). GRBs may also constrain the stellar populations in the earliest galaxies by detecting explosions from the first, metal-free stars (Population III stars; Bromm & Loeb 2006; Mészáros & Rees 2010; de Souza et al. 2011; Suwa & Ioka 2011; Toma et al. 2011; Matsumoto et al. 2015, 2016; Kinugawa et al. 2019).
A major limitation is the extreme scarcity of high-redshift ($z\geq 6$) GRB detections, with only eleven events having been identified in 20 years of Swift observations. Among these, seven have spectroscopic confirmations, while the remaining four rely on photometric redshift estimates (as listed in Table 1). This scarcity primarily results from current missions’ limited sensitivity and relatively high-energy bandpass, as high-$z$ GRBs are faint with peak emission redshifted to lower energies. Although Swift is suboptimal for detecting high-$z$ GRBs, previous studies suggest that $\sim 2\%$ of its GRBs originate at $z>5.5$ (Perley et al., 2016). However, some high-$z$ events may have been missed during the patchy follow-up campaigns. Such patchiness arises when weather conditions prevent rapid follow-up observations, when afterglows are only observed at optical wavelengths rather than in the near-infrared (NIR), and because ground facilities with deep NIR spectroscopic capabilities for high-$z$ afterglows (requiring 8 m-class telescopes) remain scarce (Campana et al., 2022).
Table 1: List of high-redshift ($z\geq 6$) GRBs detected by Swift
Spectroscopic Redshifts Photometric Redshifts
GRB $z$ Refs. GRB $z$ Refs.
050904 6.3 (1) 090429B 9.4 (12)
080913 6.7 (2) 100905A 7.9 (13)
090423 8.2 (3), (4) 120521C 6.0 (14)
130606A 5.9 (5), (6), (7) 120923A 7.8 (15)
140515A 6.3 (8), (9)
210905A 6.3 (10)
240218A 6.8 (11)
References: (1) Kawai et al. (2006); (2) Greiner et al. (2009); (3) Salvaterra et al. (2009b); (4) Tanvir et al. (2009); (5) Chornock et al. (2013); (6) Totani et al. (2014); (7) Hartoog et al. (2015); (8) Chornock et al. (2014); (9) Melandri et al. (2015); (10) Saccardi et al. (2023); (11) Saccardi et al. (2024); (12) Cucchiara et al. (2011); (13) Bolmer et al. (2018); (14) Laskar et al. (2014); (15) Tanvir et al. (2018).
To fully leverage GRBs’ potential for probing the early Universe, a much larger statistical sample of high-$z$ GRBs is essential. Consequently, detecting high-$z$ GRBs drives the design of current and future space missions like the Einstein Probe (EP; Yuan et al. 2025) and the Space-based multi-band astronomical Variable Objects Monitor (SVOM; Wei et al. 2016). GRB population studies demonstrate that optimal detection of high-$z$ GRBs requires instruments operating in the soft X-ray band with high sensitivity (Ghirlanda et al., 2015; Salvaterra, 2015).
In this work, we explore the prospects for high-$z$ GRB detection during the EP and SVOM observational era. EP was launched in January 2024, followed by SVOM in June 2024; both missions have now commenced scientific operations. Compared to other operational wide-field monitors, both the Wide-field X-ray Telescope (WXT) on board EP and the coded-mask gamma-ray imager (ECLAIRs) on board SVOM feature high sensitivity with coverage extending into relatively low X-ray energies. For a 10-s exposure, WXT achieves a sensitivity of $\sim 8.9\times 10^{-10}\,\mathrm{erg\,cm^{-2}\,s^{-1}}$ in 0.5–4 keV, while ECLAIRs reaches $\sim 7.2\times 10^{-8}\,\mathrm{erg\,cm^{-2}\,s^{-1}}$ in 4–150 keV. These properties make both instruments exceptionally well-suited for detecting high-$z$ GRBs. Notably, EP/WXT has already detected a GRB at $z=4.859$ (designated EP240315a; Liu et al. 2025), while SVOM/ECLAIRs achieved the fifth-most distant GRB redshift record with GRB 250314A at $z=7.3$ (Malesani et al., 2025), demonstrating both missions’ high-$z$ GRB survey capabilities.
This paper is structured as follows. Section 2 details our methodology for computing high-$z$ GRB detection rates for instruments with specified energy bands and sensitivities. This approach builds on a population synthesis model calibrated to Swift GRB observations. Section 3 presents the projected populations of high-$z$ GRBs detectable by EP/WXT and SVOM/ECLAIRs. We conclude with a summary and discussion in Section 4. Throughout our analysis, we adopt a standard $\Lambda$CDM cosmology with $H_{0}=67.4$ $\mathrm{km\,s^{-1}\,Mpc^{-1}}$, $\Omega_{\rm m}=0.315$, and $\Omega_{\Lambda}=0.685$ (Planck Collaboration et al., 2020).
2 Analysis Method
In this section, we describe the method to estimate the detection rate of high-$z$ GRBs for a generic detector with specified energy band and sensitivity.
2.1 The Population Model
The differential rate of GRBs detected per unit time, within redshift $z$ to $z+dz$ and luminosity $L$ to $L+dL$, is given by
$\Phi(L,\;z,\;t)=\frac{d^{3}N}{dtdzdL}=\frac{\Delta\Omega}{4\pi}\eta_{\rm duty}\frac{\psi(z)}{1+z}\frac{dV(z)}{dz}\phi(L)\;,$ (1)
where $\Delta\Omega$ is the solid angle of the sky covered by the detector, $\eta_{\rm duty}$ represents the detector’s duty cycle (i.e., the fraction of time spent performing observations and searching for GRBs) [^2] All satellites in low-Earth orbit are rendered inactive during passage through the South Atlantic Anomaly (SAA) due to elevated radiation levels. Additionally, spacecraft slewing accounts for a significant portion of the mission time (Lien et al., 2016)., $\psi(z)$ denotes the comoving formation rate of GRBs (in units of $\mathrm{Gpc^{-3}\;yr^{-1}}$), the factor $(1+z)^{-1}$ accounts for cosmological time dilation, and $\phi(L)$ is the normalized luminosity function (LF) of GRBs. Finally, the comoving volume element in a flat $\Lambda$CDM cosmology is given by
$\frac{dV(z)}{dz}=\frac{4\pi cD_{L}^{2}(z)}{(1+z)^{2}H_{0}\sqrt{\Omega_{\mathrm{m}}(1+z)^{3}+\Omega_{\Lambda}}}\;,$ (2)
where $D_{L}(z)$ is the luminosity distance at $z$.
For an instrument with a flux threshold $P_{\rm lim}$, the number of GRBs detected with redshifts within $[0,\,z_{\rm max}]$ over an observing time $T$ can be expressed as
[…]
Following Ghirlanda et al. (2015), we set the maximum redshift for our analysis to $z_{\rm max}=20$. The LF is modeled to span a range of luminosities from $L_{\rm min}=10^{47}\,\mathrm{erg\,s^{-1}}$ to $L_{\rm max}=10^{55}\,\mathrm{erg\,s^{-1}}$. For a given redshift $z$, the luminosity threshold $L_{\rm lim}$, which appears in Equation (3), is calculated by equating the peak flux of the prompt GRB emission to the detector’s sensitivity threshold $P_{\rm lim}$, i.e.,
$L_{\rm lim}(z)=4\pi D_{L}^{2}(z)P_{\rm lim}k(z)\;,$ (4)
[…]
where $N(E)$ is the observed photon spectrum (more on this below). While for other detectors with an energy flux threshold $P_{\rm lim}$ (in units of $\mathrm{erg\,cm^{-2}\,s^{-1}}$), $k(z)$ should be rewritten as
[…]
2.2 Simulation Setup
Table 2: The Properties of EP/WXT and SVOM/ECLAIRs
Mission Energy band Sensitivity (@10s@$10\,\mathrm{s}$) Field of view Duty cycle References
($\mathrm{keV}$) ($\mathrm{erg\,cm^{-2}\,s^{-1}})$ ($\mathrm{sr}$)
EP/$\mathrm{WXT}$ 0.5–4 $8.9\times 10^{-10}$ 1.1 67% Yuan et al. (2025)
SVOM/$\mathrm{ECLAIRs}$ 4–150 $7.2\times 10^{-8}$ 2.0 85% Wei et al. (2016)
The expected detection rate of GRBs for a detector with defined energy range and sensitivity becomes calculable when the intrinsic GRB formation rate, $\psi(z)$, and LF, $\phi(L)$, in Equation (3) are specified. However, these parameters remain poorly constrained. A key objective of GRB population studies is to infer them through statistical fitting of the observed distributions of redshift, luminosity, and peak flux measurements (e.g., Porciani & Madau 2001; Lloyd-Ronning et al. 2002, 2019; Firmani et al. 2004; Guetta et al. 2005; Natarajan et al. 2005; Daigne et al. 2006; Guetta & Piran 2007; Le & Dermer 2007; Salvaterra & Chincarini 2007; Kistler et al. 2008, 2009; Salvaterra et al. 2009a, 2012; Butler et al. 2010; Campisi et al. 2010; Qin et al. 2010; Wanderman & Piran 2010; Cao et al. 2011; Virgili et al. 2011; Lu et al. 2012; Robertson & Ellis 2012; Tan et al. 2013; Wei et al. 2014; Petrosian et al. 2015; Tan & Wang 2015; Yu et al. 2015; Deng et al. 2016; Pescalli et al. 2016; Wei & Wu 2017; Paul 2018; Lan et al. 2019, 2021; Le et al. 2020; Palmerio & Daigne 2021; Dong et al. 2022; Ghirlanda & Salvaterra 2022; Qu et al. 2024, 2025). Such analyses typically adopt parameterized models for both the GRB formation rate and LF, and iteratively refine them by matching simulated populations to observational data. Although the specific functional forms vary across studies, a consensus exists that GRB populations likely underwent some kind of redshift evolution (e.g., Salvaterra et al. 2009a, 2012; Virgili et al. 2011; Lan et al. 2019, 2021; Ghirlanda & Salvaterra 2022). Therefore, we employ several parameterized GRB rates and LFs, derived from recent representative population studies (Salvaterra et al., 2012; Lan et al., 2021; Ghirlanda & Salvaterra, 2022), to quantify the detection efficiency of high-$z$ GRBs for the EP/WXT and SVOM/ECLAIRs missions. Table 2 summarizes the energy bands, detection thresholds, fields of view, and duty cycles of these instruments.
In this study, we first adopt the GRB formation rate and LF derived by Salvaterra et al. (2012), hereafter S12, as our baseline framework. To ensure robustness, we also perform analogous calculations using the parameterized functions from Lan et al. (2021) and Ghirlanda & Salvaterra (2022), finding similar results across all models (see Appendix A for full comparative analysis).
Using the population synthesis framework from S12, we calculate the expected detection rate of high-$z$ GRBs by EP and SVOM. The key assumptions adopted to simulate the burst population are summarized below (for full methodological details, see Ghirlanda et al. 2015):
  1. 1.
    GRBs are distributed as a function of redshift (up to $z_{\rm max}=20$) according to the GRB formation rate (Salvaterra et al., 2012):
    $\psi(z)\propto\psi_{\star}(z)(1+z)^{\delta}\;,$ (7)
    where $\psi_{\star}(z)=(0.0157+0.118z)/[1+(z/3.23)^{4.66}]$ is the comoving cosmic star formation rate (SFR), with units of ${\rm M}_{\odot}$ $\rm yr^{-1}$ $\rm Mpc^{-3}$ (Hopkins & Beacom, 2006; Li, 2008). [^3] Note that a proportionality coefficient relating $\psi(z)$ (the GRB formation rate) to $\psi_{\star}(z)$ (the cosmic SFR) quantifies the GRB formation efficiency per unit stellar mass, with units of ${\rm M}^{-1}_{\odot}$. This coefficient is determined by normalizing the model predictions to the observed population of GRBs. S12 found that a strong redshift-dependent evolution in the GRB rate density ($\delta=1.7\pm 0.5$) is required to explain the observations.
  2. 2.
    The GRB LF is assumed as a broken power law (Salvaterra et al., 2012):
    $\phi(L)\propto\left\{\begin{array}[]{ll}\left(\frac{L}{L_{b}}\right)^{-\nu_{1}};~{}~{}~{}~{}~{}~{}~{}L\leq L_{b}\\ \left(\frac{L}{L_{b}}\right)^{-\nu_{2}};~{}~{}~{}~{}~{}~{}~{}L>L_{b}\;,\\ \end{array}\right.$ (8)
    where $\nu_{1}$ and $\nu_{2}$ are the power-law indices describing the LF below and above the break luminosity $L_{b}$, respectively. S12 derived the best-fit parameters $\nu_{1}=1.50^{+0.16}_{-0.32}$, $\nu_{2}=2.32^{+0.77}_{-0.32}$, and $L_{b}=(3.8^{+6.3}_{-2.7})\times 10^{52}\,\mathrm{erg\,s^{-1}}$. Here, the GRB LF is normalized to unity by integrating over the luminosity range $L_{\rm min}=10^{47}\,\mathrm{erg\,s^{-1}}$ to $L_{\rm max}=10^{55}\,\mathrm{erg\,s^{-1}}$.
  3. 3.
    To compute the peak flux $P$ of GRBs in a given energy band, it is necessary to associate a spectral model $N(E)$ to each mock burst characterized by redshift $z$ and luminosity $L$ (see Equations (4)–(6)). Here we adopt a Band function spectrum (Band et al., 1993), with the low-energy photon spectral index $\alpha$ and high-energy photon spectral index $\beta$ sampled from Gaussian distributions centered at $-1$ and $-2.3$, respectively, both with a standard deviation of $\sigma=0.2$ (Kaneko et al., 2006; Nava et al., 2011; von Kienlin et al., 2020).
  4. 4.
[…]
2.3 Simulation Data Calibration and Consistency Verification
Our simulated GRB population is calibrated by the observed population of bright bursts detected by the Burst Alert Telescope (BAT) on board the Swift satellite. S12 constructed a complete Swift/BAT sample (the BAT6 sample), consisting of 58 GRBs selected under criteria favorable for redshift measurement (Jakobsson et al., 2006) and with peak flux $P\geq 2.6$ $\mathrm{\rm ph\,cm^{-2}\,s^{-1}}$ in the 15–150 keV energy range. Applying the same flux threshold used to define the BAT6 sample, we select 453 GRBs with 1-s peak flux $P\geq 2.6$ $\mathrm{\rm ph\,cm^{-2}\,s^{-1}}$ from the most recent Swift sample (1425 GRBs with duration $T_{90}\geq 2\,\mathrm{s}$, detected up to December 2024). [^4] https://swift.gsfc.nasa.gov/archive/grb_table/ Considering Swift/BAT’s field of view ($\mathrm{1.4\,sr}$), a mission period of $\sim 20\,\mathrm{years}$, and an average instrument duty cycle of 78% (Lien et al., 2016), the detection rate of GRBs with peak flux $P\geq 2.6$ $\mathrm{\rm ph\,cm^{-2}\,s^{-1}}$ (15–150 keV) is estimated to be $\sim 21$ $\mathrm{events\,yr^{-1}\,sr^{-1}}$. Our simulated GRB population is scaled to match this observed rate.
Figure 1: Comparison of cumulative flux distributions between the simulated GRB population (solid lines) and observed burst samples (dots). Blue and yellow dots represent the observed distributions for the Swift/BAT (453 GRBs; $P\geq 2.6$ $\mathrm{\rm ph\,cm^{-2}\,s^{-1}}$) and Fermi/GBM (2637 GRBs; $P\geq 0.7$ $\mathrm{\rm ph\,cm^{-2}\,s^{-1}}$) samples, respectively. Swift/BAT fluxes are integrated in the 15–150 keV energy range, while Fermi/GBM fluxes use 50–300 keV. Solid lines (color-matched to observations) show the simulated GRB population’s flux distributions for each instrument’s energy range. No attempt was made to fit the observed flux distributions.
Figure 1 displays the cumulative flux distribution of the observed Swift sample (blue dots), comprising 453 GRBs with $P\geq 2.6$ $\mathrm{\rm ph\,cm^{-2}\,s^{-1}}$. Using the simulation framework described above, we compute the expected flux distribution for Swift/BAT by adopting a detection threshold of $P_{\rm lim}=2.6$ $\mathrm{\rm ph\,cm^{-2}\,s^{-1}}$ in the 15–150 keV energy band. The simulated GRB population’s flux distribution is shown as the blue solid line in Figure 1. We find that the expected flux distribution matches the observed Swift sample well without requiring adjustment to the model parameters. This consistency supports the conclusion that the model assumptions (Section 2.2), on which the code is based, are correctly implemented and successfully reproduce the population of Swift bursts from which the model is derived.
To determine whether the simulated GRB population is also representative of samples observed by instruments beyond Swift/BAT, we compare the cumulative flux distribution of the simulated population with that of the observed Fermi sample. The public catalog of GRBs detected by Fermi’s Gamma-ray Burst Monitor (Fermi/GBM) includes 3249 bursts with $T_{90}\geq 2\,\mathrm{s}$, recorded between July 2008 and December 2024 (Gruber et al., 2014; von Kienlin et al., 2014; Narayana Bhat et al., 2016; von Kienlin et al., 2020). [^5] https://heasarc.gsfc.nasa.gov/W3Browse/fermi/fermigbrst.html To address potential incompleteness in faint burst sampling, we consider the flux integrated over the 50–300 keV energy range and apply a cut to the GBM sample at $P\geq 0.7$ $\mathrm{\rm ph\,cm^{-2}\,s^{-1}}$, resulting in 2637 GBM bursts. Considering an average sky coverage fraction of $\sim 70\%$ for GBM, a mission duration of $\sim 16.5\,\mathrm{years}$ (encompassing our sample period), and an average instrument duty cycle of 85% (von Kienlin et al., 2020), the GBM detection rate for bursts with $P\geq 0.7$ $\mathrm{\rm ph\,cm^{-2}\,s^{-1}}$ in the 50–300 keV energy range is $\sim 21$ $\mathrm{events\,yr^{-1}\,sr^{-1}}$. Figure 1 shows that the cumulative flux distribution of the 2637 observed Fermi GRBs (yellow dots) aligns well with the corresponding distribution of the simulated GRB population (yellow solid line). Again, this agreement reinforces the reliability of our analysis.
3 Populations of High-z GRBs Accessible by EP and SVOM
Using the calibrated synthetic GRB population from S12, we can predict the detection rate of an instrument operating in a specific energy band once its flux threshold is known. Figure 2 presents the redshift distributions and cumulative numbers of detected GRBs expected from observations by EP/WXT and SVOM/ECLAIRs (red curves for EP/WXT and blue ones for SVOM/ECLAIRs). The shaded regions around the model line represent the 68% uncertainty in the model predictions. These uncertainties are estimated through 1000 Monte Carlo simulations incorporating uncertainties in the model parameters.
For EP/WXT, we adopt a field of view of $\mathrm{1.1\,sr}$, a duty cycle of 67%, and a flux sensitivity of $8.9\times 10^{-10}$ $\mathrm{\rm erg\,cm^{-2}\,s^{-1}}$ for a 10-s exposure in the 0.5–4 keV energy range (Yuan et al., 2025). Within the models of S12, the total number of GRB detections expected for a 1-year operation is $\simeq 276^{+264}_{-175}$ ($1\sigma$ confidence intervals; red curve in the bottom panel of Figure 2), which may exceed the actual observed rate. As of May 2025, $\simeq 90$ fast X-ray transients have been reported in General Coordinated Network (GCN) circulars, [^6] https://gcn.nasa.gov/circulars corresponding to $\simeq 68$ $\mathrm{events\,yr^{-1}}$. However, not all detections are currently reported, and instrument characteristics remain under investigation. In addition, our choice of the limiting flux, $\sim 8.9\times 10^{-10}$ $\mathrm{\rm erg\,cm^{-2}\,s^{-1}}$, may not be appropriate for comparing our estimates with the current observed detection count. This is because faint GRBs with peak fluxes slightly above the EP/WXT threshold are not always reliably detected. Under the 0.5–4 keV sensitivity parameters of WXT, our implementation of the S12 framework yields high-$z$ GRB detection rates of $\sim 5.1^{+3.4}_{-2.4}$ ($z>6$), $\sim 1.3^{+1.2}_{-0.7}$ ($z>8$), and $\sim 0.5^{+0.5}_{-0.3}$ ($z>10$) $\mathrm{events\,yr^{-1}}$, with uncertainties corresponding to $1\sigma$ confidence levels. These rates are listed in Columns 3-5 of Table 3.
We implement SVOM/ECLAIRs including a field of view of $\mathrm{2.0\,sr}$, 85% duty cycle, and 4–150 keV sensitivity threshold of $7.2\times 10^{-8}$ $\mathrm{\rm erg\,cm^{-2}\,s^{-1}}$ for a 10-s exposure (Wei et al., 2016). Based on the models by S12, the predicted total number for a 1-year operation is $\simeq 83^{+16}_{-15}$ ($1\sigma$ confidence intervals; blue curve in the bottom panel of Figure 2), roughly consistent with the observed rate. As of May 2025, $\simeq 40$ GRBs have been detected by SVOM/ECLAIRs (reported in GCN circulars), corresponding to $\simeq 48$ $\mathrm{events\,yr^{-1}}$. We estimate that the detection rates of high-$z$ GRBs observable by SVOM/ECLAIRs are $\sim 0.7^{+1.0}_{-0.4}$, $\sim 0.1^{+0.3}_{-0.1}$, and $\sim 0.04^{+0.08}_{-0.03}$ $\mathrm{events\,yr^{-1}}$ at $z>6$, $z>8$, and $z>10$, respectively ($1\sigma$ confidence levels; see Columns 7-9 of Table 3).
Figure 2: Top panel: Redshift distributions of GRBs detectable by EP/WXT (red curve) and SVOM/ECLAIRs (blue curve), derived from the population synthesis framework of S12. Bottom panel: Cumulative detection rates for EP/WXT and SVOM/ECLAIRs, with color coding consistent with the top panel. Shaded regions, estimated by 1000 Monte Carlo simulations, represent $1\sigma$ confidence intervals incorporating uncertainties in the model parameters.
Table 3: Cumulative Detection Rates of GRBs by EP/WXT and SVOM/ECLAIRs Across Population-Synthesis Models
Detection Rate $N_{\rm GRB}(>z)$ $(\mathrm{events\,yr^{-1}})$
Model EP/WXT SVOM/ECLAIRs
$z>0$ $z>6$ $z>8$ $z>10$ $z>0$ $z>6$ $z>8$ $z>10$
Salvaterra et al. (2012) $276^{+264}_{-175}$ $5.1^{+3.4}_{-2.4}$ $1.3^{+1.2}_{-0.7}$ $0.5^{+0.5}_{-0.3}$ $83^{+16}_{-15}$ $0.7^{+1.0}_{-0.4}$ $0.1^{+0.3}_{-0.1}$ $0.04^{+0.08}_{-0.03}$
Lan et al. (2021) $343^{+69}_{-54}$ $3.5^{+0.9}_{-0.7}$ $0.9^{+0.3}_{-0.2}$ $0.3^{+0.1}_{-0.1}$ $80^{+1}_{-1}$ $0.7^{+0.2}_{-0.2}$ $0.1^{+0.1}_{-0.1}$ $0.02^{+0.02}_{-0.01}$
Ghirlanda & Salvaterra (2022) $84^{+29}_{-24}$ $4.7^{+4.3}_{-2.3}$ $1.7^{+2.0}_{-1.0}$ $0.7^{+1.0}_{-0.4}$ $74^{+10}_{-11}$ $1.3^{+2.1}_{-0.8}$ $0.3^{+0.8}_{-0.2}$ $0.11^{+0.31}_{-0.08}$
Figure 3: Peak isotropic luminosity versus redshift of the simulated GRB populations detectable by EP/WXT (red solid contours) and SVOM/ECLAIRs (blue solid contours), with sensitivity thresholds for a 10-s exposure. The shaded regions represent the $1-3\sigma$ confidence contours. The blue dashed contours correspond to SVOM/ECLAIRs with a sensitivity threshold for a 1000-s exposure. The eleven high-redshift ($z\geq 6$) GRBs detected by Swift are indicated by star symbols. A square symbol marks the detection of long GRB 250314A at $z=7.3$ by SVOM/ECLAIRs.
By incorporating the properties of EP/WXT and SVOM/ECLAIRs (as listed in Table 2), we run our population synthesis code to perform Monte Carlo simulations. Figure 3 displays the contour levels in the plane of the peak isotropic luminosity $L$ versus redshift $z$ for the simulated GRB populations detectable by EP/WXT (red solid contours) and SVOM/ECLAIRs (blue solid contours). For comparison, we include the eleven high-redshift ($z\geq 6$) GRBs detected by Swift (star symbols). [^7] The relevant information for the peak luminosities of these eleven Swift GRBs is available from published literature. For GRB 210905A, we used $L=1.87\times 10^{53}$ $\mathrm{erg\,s^{-1}}$ from Rossi et al. (2022). For GRB 240218A, we calculated $L=2.88\times 10^{53}$ $\mathrm{erg\,s^{-1}}$ using the peak photon flux and spectral parameters reported in Veres et al. (2024). For the other nine GRBs, peak luminosities were taken from Lan et al. (2021). Due to its extended energy range into soft X-rays and higher sensitivity, the EP/WXT simulations sample relatively less energetic events compared to those currently detected by Swift and those expected from SVOM in the near future. GRB 250314A at $z=7.3$, [^8] The peak luminosity of GRB 250314A is $L=1.97\times 10^{52}$ $\mathrm{erg\,s^{-1}}$, based on private communication with the SVOM team. detected by SVOM/ECLAIRs (square symbol; Malesani et al. 2025), lies outside the $3\sigma$ confidence region predicted by SVOM/ECLAIRs simulations (blue solid contours). In our simulations, the fiducial sensitivity of SVOM/ECLAIRs is set to $7.2\times 10^{-8}$ $\mathrm{erg\,cm^{-2}\,s^{-1}}$, achievable with an exposure time of $\simeq 10\,\mathrm{s}$. The instrument’s maximal sensitivity, however, reaches $7.2\times 10^{-10}$ $\mathrm{\rm erg\,cm^{-2}\,s^{-1}}$ for exposures of $\simeq 1000\,\mathrm{s}$ (Wei et al., 2016). Adopting this higher sensitivity threshold, the simulated GRB population detectable by SVOM/ECLAIRs shifts to the blue dashed contours. Notably, GRB 250314A now falls within the $3\sigma$ confidence region of these updated predictions. In the 15–150 keV energy band, Swift-detected long GRBs typically last tens to hundreds of seconds (Horváth & Tóth, 2016). At lower energies ($\mathrm{<10\,keV}$), they show substantially longer durations, with EP-observed GRBs averaging several hundred seconds (Gao et al., 2025). While cosmological time-dilation would prolong observed durations at high redshifts, bandpass-shifting and undetected weak prompt emission episodes counteract this effect (Littlejohns et al., 2013). These factors collectively justify longer exposures for soft X/gamma-ray instruments like EP/WXT and SVOM/ECLAIRs.
Figure 4: Peak isotropic luminosity versus rest-frame peak energy of the simulated GRB populations detectable by EP/WXT (red solid contours) and SVOM/ECLAIRs (blue solid contours), with detection sensitivity thresholds corresponding to a 10-s exposure. The shaded regions represent the $1-3\sigma$ confidence contours. Yellow points indicate real GRBs detected by Swift (adopted from Pescalli et al. 2016).
Figure 4 compares the peak isotropic luminosity ($L$) versus rest-frame peak energy ($E_{\rm peak}(1+z)$) of Swift-detected GRBs (yellow dots) with simulated GRB populations detectable by EP/WXT (red solid contours) and SVOM/ECLAIRs (blue solid contours). The softer energy band and enhanced sensitivity of EP/WXT enable it to detect a distinct GRB population dominated by soft spectra and low luminosities (primarily low-z events), while simultaneously detecting high–z GRBs whose spectra are softened by cosmological redshift effects. Compared to Swift and SVOM, EP/WXT will advance the study of low-luminosity GRBs, which may exhibit different physical properties (e.g., opening angles or jet bulk velocities). Current knowledge of this population is limited to a few bursts due to the higher energy bands covered by existing missions. EP/WXT’s unique capability to trigger in the X-ray band will open a new window for studying low-$z$, soft-spectrum GRBs, which are predicted to dominate the total GRB population. Based on our population models, we estimate that $\sim 10\%$ of bursts detected at $z<1$ will have luminosities less than $\mathrm{10^{49}\;erg\,s^{-1}}$ and soft spectra with peak energies $<20\,\mathrm{keV}$. A statistically robust sample of these soft, low-luminosity GRBs will provide critical insights into the nature of X-ray flashes, X-ray rich events (Sakamoto et al., 2008), and low-luminosity GRBs (Liang et al., 2007; Pescalli et al., 2015; Salafia et al., 2016), whose origins remain unclear. Finally, the ratio of low-luminosity GRBs to supernovae could constrain the efficiency with which massive stars produce relativistic jets capable of breaking out of stellar envelopes (Ghirlanda et al., 2013).
4 Summary and Discussions
In this work, we predict high-$z$ GRB detection rates for EP/WXT and SVOM/ECLAIRs. WXT’s soft X-ray band (0.5–4 keV) and exceptional sensitivity ($\sim 8.9\times 10^{-10}\,\mathrm{erg\,cm^{-2}\,s^{-1}}$) make it ideally suited for identifying high-$z$ GRBs, as demonstrated by its detection of EP240315a at $z=4.859$ (Liu et al., 2025). ECLAIRs’ 4 keV low energy threshold similarly facilitates high-$z$ detections, exemplified by GRB 250314A at $z=7.3$ (Malesani et al., 2025). Based on a physically motivated population synthesis model calibrated to Swift observations, we evaluate the redshift distributions and detection rates of GRBs for both instruments. Our results indicate that EP/WXT could detect $\sim 5.1^{+3.4}_{-2.4}$ GRBs at $z>6$ for a 1 yr operation, while SVOM/ECLAIRs is expected to detect $\sim 0.7^{+1.0}_{-0.4}$ $\mathrm{events\,yr^{-1}}$ at $z>6$. In addition to the representative GRB population model of S12, we perform analogous calculations using alternative population synthesis models from Lan et al. (2021) and Ghirlanda & Salvaterra (2022). We derive consistent high-$z$ detection rates across all models, demonstrating robustness to population synthesis assumptions.
Note that EP, operating solely as an X-ray telescope, cannot determine redshifts. Optical/NIR follow-up observations are essential for identifying high-$z$ GRBs, though such follow-ups may not always be feasible for EP-detected events. This could yield fewer confirmed high-$z$ GRBs than our estimates suggest. Assuming a $\sim 30\%$ efficiency for ground-based spectroscopic redshift determination, we expect EP to detect $\sim 1.5^{+1.0}_{-0.7}$ GRBs at $z>6$ annually. Equipped with dedicated follow-up telescopes, SVOM will play a crucial role in promptly identifying high-$z$ GRB candidates deserving deep NIR spectroscopy (see Llamas Lanza et al. 2024). Specifically, very high-redshift cases may be indicated by non-detection in both channels of the on-board Visible Telescope (VT) (though not uniquely, as dusty GRBs would also appear optically faint). These candidates can then be prioritized for ground-based NIR follow-up observations using positional locations from SVOM’s Microchannel X-ray Telescope (MXT).
After several years of operation, we anticipate that EP and SVOM will double the number of existing high-$z$ GRBs. Furthermore, their substantial sample of low-luminosity GRBs at low-to-intermediate redshifts will enable investigations into potential sub-populations with dominant soft X-ray flashes and the faint end of the GRB LF—both offering key insights into GRB jet structure (Liang et al., 2007; Pescalli et al., 2015; Salafia et al., 2016). In conclusions, EP and SVOM will open new avenues for advancing GRB studies and early Universe exploration through high-$z$ beacons that probe the cosmic frontier. Notably, besides EP and SVOM, several mission concepts specifically designed for high-$z$ GRB discovery, including THESEUS (Amati et al., 2018, 2021; Ghirlanda et al., 2021), Gamow Explorer (White, 2020), and HiZ-GUNDAM (Yonetoku et al., 2024), are under active investigation.
Acknowledgments
We would like to thank the anonymous referee for valuable comments that helped improve our manuscript. We are also grateful to Giancarlo Ghirlanda, Hou-Dun Zeng, Guang-Xuan Lan, Chong-Yu Gao, Chen-Wei Wang, and Wen-Long Zhang for useful discussions. This work is supported by the Strategic Priority Research Program of the Chinese Academy of Sciences (grant No. XDB0550400), the National Key R&D Program of China (2024YFA1611704), the National Natural Science Foundation of China (grant Nos. 12422307, 12373053, and 12321003), the Key Research Program of Frontier Sciences (grant No. ZDBS-LY-7014) of Chinese Academy of Sciences, and the Natural Science Foundation of Jiangsu Province (grant No. BK20221562).
Appendix A Alternative GRB Formation Rate and LF
We briefly discuss results for alternative GRB formation rates and LFs. Alongside the representative GRB population model of S12, we consider analyses from Lan et al. (2021) and Ghirlanda & Salvaterra (2022).
Figure 5: Same as Figure 2, but now for the GRB formation rate and LF models of Lan et al. (2021) and Ghirlanda & Salvaterra (2022).
Lan et al. (2021) analyzed 302 GRBs detected by Swift up to November 2019 with peak flux $P\geq 1.0$ $\mathrm{\rm ph\,cm^{-2}\,s^{-1}}$ in the 15–150 keV energy range. They considered three different scenarios: (i) both a non-evolving GRB formation rate (where the GRB formation rate strictly follows the cosmic SFR) and a non-evolving LF; (ii) a redshift-evolving GRB formation rate with a non-evolving LF; and (iii) a non-evolving GRB formation rate with a redshift-evolving LF. Here we adopt their results for scenario (ii). That is, the GRB formation rate is proportional to the SFR with an evolutionary term parameterized by $(1+z)^{\delta}$, as in Equation (7), and the non-evolving LF is described by a broken power law:
$\phi(L)=\frac{A}{\ln(10)L}\left\{\begin{array}[]{ll}\left(\frac{L}{L_{b}}\right)^{-\nu_{1}};~{}~{}~{}~{}~{}~{}~{}L\leq L_{b}\\ \left(\frac{L}{L_{b}}\right)^{-\nu_{2}};~{}~{}~{}~{}~{}~{}~{}L>L_{b}\;,\\ \end{array}\right.$ (A1)
where $A$ is a normalization constant. [^9] To maintain consistency with the S12 model, $A$ is determined by normalizing the LF such that its integral equals unity over the luminosity range $L_{\rm min}=10^{47}\,\mathrm{erg\,s^{-1}}$ to $L_{\rm max}=10^{55}\,\mathrm{erg\,s^{-1}}$. The model parameters are obtained as $\delta=1.43^{+0.22}_{-0.20}$, $\nu_{1}=0.60^{+0.05}_{-0.05}$, $\nu_{2}=1.65^{+0.27}_{-0.28}$, and $\log_{10}\left(L_{b}/\,\mathrm{erg\,s^{-1}}\right)=52.98^{+0.11}_{-0.12}$. We extracted 942 GRBs with peak flux $P\geq 1.0$ $\mathrm{\rm ph\,cm^{-2}\,s^{-1}}$ (i.e., the same flux threshold used to define the sample of Lan et al. (2021)) from the most recent Swift sample of 1425 GRBs (all with $T_{90}\geq 2\,\mathrm{s}$) detected through December 2024. Considering a field of view of $\mathrm{1.4\,sr}$, a mission period of $\sim 20\,\mathrm{years}$, and an average instrument duty cycle of 78%, the Swift/BAT detection rate for bursts with peak flux $P\geq 1.0$ $\mathrm{\rm ph\,cm^{-2}\,s^{-1}}$ in the 15–150 keV band is $\sim 43$ $\mathrm{events\,yr^{-1}\,sr^{-1}}$. Within the population synthesis framework of Lan et al. (2021), we calibrate our simulated GRB population to this observed rate.
Ghirlanda & Salvaterra (2022) analyzed the BAT6 sample, which consists of GRBs with a 15–150 keV peak flux $P\geq 2.6$ $\mathrm{\rm ph\,cm^{-2}\,s^{-1}}$. They parameterized the GRB formation rate as:
$\psi(z)\propto\frac{(1+z)^{q_{1}}}{1+\left(\frac{1+z}{q_{2}}\right)^{q_{3}}}\;,$ (A2)
a functional form motivated by its established use for fitting the cosmic SFR (Madau & Dickinson, 2014). Their LF adopts a broken power-law form defined for $L\geq 10^{47}\,\mathrm{erg\,s^{-1}}$ with a redshift-evolving break luminosity:
$\phi(L,\,z)\propto\left\{\begin{array}[]{ll}\left(\frac{L}{L_{b,z}}\right)^{-\nu_{1}};~{}~{}~{}~{}~{}~{}~{}L\leq L_{b,z}\\ \left(\frac{L}{L_{b,z}}\right)^{-\nu_{2}};~{}~{}~{}~{}~{}~{}~{}L>L_{b,z}\;,\\ \end{array}\right.$ (A3)
where $L_{b,z}=L_{b,0}(1+z)^{\delta}$. The model parameters are obtained as $q_{1}=3.33^{+0.33}_{-0.33}$, $q_{2}=3.42^{+0.28}_{-0.28}$, $q_{3}=6.21^{+0.38}_{-0.32}$, $\nu_{1}=0.97^{+0.05}_{-0.04}$, $\nu_{2}=2.21^{+0.13}_{-0.18}$, $\log_{10}\left(L_{b,0}/\,\mathrm{erg\,s^{-1}}\right)=52.02^{+0.22}_{-0.19}$, and $\delta=0.64^{+0.32}_{-0.26}$. Within the population synthesis framework of Ghirlanda & Salvaterra (2022), we adopt the same Swift/BAT detection rate ($\sim 21$ $\mathrm{events\,yr^{-1}\,sr^{-1}}$) for bursts with peak flux $P\geq 2.6$ $\mathrm{\rm ph\,cm^{-2}\,s^{-1}}$ in the 15–150 keV band as used in the S12 model to calibrate our simulated GRB population.
Figure 5 shows the redshift distributions and cumulative numbers of GRBs detectable by EP/WXT and SVOM/ECLAIRs under the Lan et al. (2021) and Ghirlanda & Salvaterra (2022) models. The expected GRB detection rates at various redshifts by EP/WXT and SVOM/ECLAIRs for these models are summarized in Table 3. Although the adopted GRB formation rates and LFs differ from those in the S12 model, the high-$z$ GRB detection rates show qualitative similarities.